Circular No. 8150

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003gc Further to IAUC 8149, W. M. Wood-Vasey, G. Aldering, and P. Nugent

report the discovery of an apparent supernova (mag 18.0) on unfiltered NEAT images taken on Mar. 27.14–27.18 UT. The new object is located at α = 10^{h}54^{m}22^{s}.26, δ = −6^{o}15′48′′.9 (equinox 2000.0), which is 1′′.0 west and 0′′.4 north of the center of the host galaxy. Additional NEAT magnitudes: 2002 Apr. 8.15, [20.7; May 11.19, [20.6; 2003 Feb. 23.53, 17.2; Mar. 28.13, 17.9; Apr. 1.24, 18.0; May 6.24, 19.0.

SUPERNOVA 2003gd IN M74 R. Evans, Hazelbrook, N.S.W., reports his visual discovery with a 0.31-

m reflector of an apparent supernova (mag ∼ 13.2) on June 12.82 UT, located on top of the southern spiral arm of the galaxy, 20′′ east and 150′′ south of the center of M74 = NGC 628. R. H. McNaught, Australian National University, reports confirmation of the new object (red mag 13.9) on images taken on June 13.84 with the 1.0-m telescope at Siding Spring Observatory, with a precise position measured as α = 1^{h}36^{m}42^{s}.65, δ = +15^{o}44′20′′.9 (equinox 2000.0; poor fit with ∼ 50 USNO-A2.0 stars; uncertainties ± 0′′.6 in α, ± 0′′.3 in δ), which is 13′′.2 east and 161′′ south of the center of M74. A foreground star of red mag 12.4 has position end figures 32^{s}.00, 45′08′′.7. McNaught adds that SN 2003gd is absent from Digitized Sky Survey images.

P. Garnavich, University of Notre Dame; and E. Bass, Cornell University, obtained near-infrared spectra (range 850–2400 nm) of SN 2003gd in twilight on June 13.46 UT with the 1.8-m Vatican Advanced Technology Telescope (+ CorMASS spectrograph; low-dispersion echelle design). Preliminary reduction of the J-band order shows strong, broad emission of Paβ and Paγ, which confirms that this is a supernova and indicates it to be a type-II event.

SUPERNOVA 2002lg Further to IAUC 8141, L.-G. Strolger et al. also report an appar-

ent supernova (z = 23.4) from ACS z-band images of the ‘Chandra Deep Field South’ taken on 2002 July 4.2 UT. SN 2002lg was located at α = 3^{h}32^{m}35^{s}.76, δ = −27^{o}47′58′′.8 (equinox 2000.0), which is 0′′.13 east and 0′′.10 south of the center of its apparent host galaxy. Additional z magnitudes for SN 2002lg: 2002 Aug. 2.9, 26.0; Sept. 19.8, [26.5.

2003 June 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green